OpenStudy (anonymous):

A large diamond with a mass of 481.3 grams was recently discovered in a mine. If the density of the diamond is 3.51 grams over centimeters cubed, what is the volume? Round your answer to the nearest hundredth. A. 48.23 cm3 B. 55.91 cm3 C. 58.71 cm3 D.137.1 cm3

OpenStudy (anonymous):

@mathslover

mathslover (mathslover):

What you are given with in the question?

OpenStudy (anonymous):

The mass of grams and the density.

mathslover (mathslover):

Not mass of grams but Mass of the large diamond (in grams) and the density of diamond.

OpenStudy (anonymous):

So, do I divide the two?

mathslover (mathslover):

You know Density = Mass/ Volume 3.51 g/cm^3 = 481.3 g / Volume Volume = 481.3 g / 3.51 g/cm^3 = \(\dfrac{481.3 }{3.51}\) cubic cm .

OpenStudy (anonymous):

I got it. It was D, if I was right.

mathslover (mathslover):

Yep, you're right... No wait.. you're right \(\LARGE{\mathsf{AGAIN}}\) ! :)